how many grams of 4% salt solution must be mixed with 50g of a 9% salt solution to make a 6% salt solution?
Answer by stanbon(75887)

how many grams of 4% salt solution must
be mixed with 50g of a 9% salt solution
to make a 6% salt solution?
--------------
Equation:
salt + salt = salt
0.04x + 0.09*50 = 0.06(x+50)
--------
Multiply thru 100 to get:
4x + 9*50 = 6x + 6*50
2x = 3*50
x = 75 grams (amt. of 4% solution needed for the mixture)
===============
Cheers,
Stan H.
===============